Find the measure of an angle such that the difference between the measure
of its supplement and 3 times the measure of the complement is 16 degrees.
: Find the measure of an angle such that the difference between the measure
of its supplement and 3 times the measure of the complement is 16 degrees.

Answer by edjones(2415)

Let x=the angle
(180-x)-3(90-x)=16
180-x-270-3x=16
-4x=16-180-270
-4x=-434
x=108.5 deg.
.
Ed